OpenStudy (anonymous):

if elevator goes up at 3.125 m/s2 for 4.0 seconds what is a. final speed in meters per second ? b. final speed in miles per hour ?

OpenStudy (anonymous):

a. 3.125 * 4 =12.5m/s if i'm not mistaken

OpenStudy (anonymous):

^^ that's right assuming zero initial velocity.

OpenStudy (anonymous):

b. \[12\times1000\div60\times1\div0.62\times1\div60\] =5.378mile/hour is it anwar? o.62 = 1km 1km=1000m

OpenStudy (anonymous):

since we have the following formula: a)\[\nu=\nu_i+at \implies v=0+(3.125)(4)=12.5 m/s\] b)\[1 meter = 0.62\times10^{-3} miles, 1 hour =3600 \sec\] the rest is trivial

OpenStudy (anonymous):

anwar,why dont u help me solve my differential equation problem..... ^^" if that doesnt bother u

OpenStudy (anonymous):

I hope it's an easy one, because it's already my bed time :(

OpenStudy (anonymous):

i've already post it